# Constrained Stochastic Recursive Momentum Successive Convex Approximation

### v2

@@ IV Complexity analysis
The SFO complexity bound of Theorem 2 matches with that of STORM [33], STORM+ [27], and nearly matches with that of Pstorm [30], where the complexity is $\mathcal{O}(\epsilon^{-3/2})$ for the unconstrained case. Remarkably, while STORM and Pstorm were proposed for unconstrained and convex constrained problems respectively, CoSTA can also handle non-convex constraints at almost similar SFO complexity. It is remarked that the non-asymptotic results in Theorem 2 can also be used to infer the asymptotic convergence of the corresponding quantities.
@@ V-A Sparse Binary Classification
The problem of finding a sparse binary classifier can be written in the form of ($\mathcal{P^{\prime}}$) with the objective being the logistic loss function, i.e., $f({\mathbf{x}},\{{\mathbf{a}},b\})=\log(1+e^{-b{\mathbf{a}}^{\mathsf{T}}{\mathbf{x}}})$ and the constraint $g({\mathbf{x}})$ taking form of a sparsity inducing function, such as the $\ell_{1}$ norm or the minimax concave function [8]. The random variable here signifies the independent observations $\{{\mathbf{a}},b\}$.
For our case we will take $g$ to be minmax concave penalty (MCP) [8], given by:
$g({\mathbf{x}})$ $=\lambda\|{\mathbf{x}}\|_{1}-\sum_{k=1}^{n}h_{\lambda,\theta}([{\mathbf{x}}]_{k})$ (60)
where $[{\mathbf{x}}]_{k}$ denotes the $k$-th entry of ${\mathbf{x}}$ and $h_{\lambda,\theta}$ is defined as:
$h_{\lambda,\theta}(x)=\begin{cases}\frac{x^{2}}{2\theta},&|x|\leq\theta\lambda\\ \lambda|x|-\frac{\theta\lambda^{2}}{2},&|x|>\theta\lambda.\\ \end{cases}$ (61)
@@
We can obtain a smoothed version of the MCP by approximating $|x|$ as $\sqrt{x^{2}+\varrho}$ where $\varrho$ is a positive constant. Since the objective is already convex, its running surrogate ${\tilde{f}}$ can be constructed using ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ and (6) as
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})$ (62)
@@
$+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$
where ${\boldsymbol{\xi}}_{t}=\{{\mathbf{a}}_{t},b_{t}\}$. On the other hand, the surrogate of $h_{\lambda,\theta}$ can be constructed as,
$h_{\lambda,\theta}(x,x_{t})$ $=h_{\lambda,\theta}(x_{t})+\nabla h_{\lambda,\theta}(x_{t})(x-x_{t})$ (63)
=$=\begin{cases}\frac{x_{t}x}{\theta}-\frac{x_{t}^{2}}{2\theta},&|x_{t}|\leq\theta\lambda\\ \frac{\lambda\varrho}{\sqrt{x_{t}^{2}+\varrho}}-\frac{\theta\lambda^{2}}{2}+\frac{\lambda x_{t}x}{\sqrt{x_{t}^{2}+\varrho}},&|x_{t}|>\theta\lambda.\\ \end{cases}$ (64)
Hence, the surrogate of the constraint (60) at point ${\mathbf{x}}_{t}$ becomes
@@
${\tilde{g}}({\mathbf{x}},{\mathbf{x}}_{t})=\sum_{k=1}^{n}\left(\lambda\sqrt{[{\mathbf{x}}]^{2}_{k}+\varrho}-h_{\lambda,\theta}([{\mathbf{x}}]_{k},[{\mathbf{x}}_{t}]_{k})\right)$ (65)
@@
We compare the performance of our algorithm with that of the LCPP method, which was proposed in [8] for solving stochastic non-convex optimization problems with sparsity inducing non-convex constraints. It is proximal point algorithm that solves the a sequence of convex subproblems with gradually relaxed constraint levels.
@@
For our simulations use two datasets: MNIST and gisette, details of which are summarized in Table II. For the MNIST dataset, we consider the binary classification problem of classifying the handwritten digit 5 from the other digits. As the dataset consists of 10 classes this can be achieved by assigning a label $+1$ to digit $5$ and $-1$ to other digits. Likewise, gisette dataset also consists handwritten digits, but here the classification is between the often confused digits 4 and 9.
@@
TABLE IV: Average time per iteration in seconds for $\sigma$ = 0.2, $Tf$ = 15, environmental configuration 4, and different values of $T$.
@@ VI Conclusion
We have considered the problem of minimizing a stochastic non-convex optimization problem with non-convex constraints. While the problem has been well-studied, the state-of-the-art algorithms have suboptimal oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$. We have proposed the CoSTA algorithm that introduces recursive momentum updates, thereby achieving the unimprovable oracle complexity of ${\mathcal{O}}(\epsilon^{-3/2})$. The proposed algorithm is developed within the SCA framework, and requires convex surrogates of the objective and constraint functions at every iteration. We also introduced a parameterized version of the Mangasarian-Fromovitz constraint qualification that allows us to ensure that the intermediate subproblems are strictly feasible and the intermediate dual variable stay bounded, resulting in an oracle complexity bound that depends only on the problem parameters. Finally, the empirical performance of the proposed CoSTA algorithm is compared with the other state-of-the-art algorithms across two different applications: energy optimal trajectory design and sparse binary classification. The proposed algorithm outperforms the state-of-the-art algorithms on the energy optimal trajectory design problem and performs at par with the LCPP algorithm on the sparse binary classification problem.
@@ Appendix A Proof of Lemma 1
The proof follows along the lines of [33, Lemma 2]. Expanding ${\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})$ from (7) and introducing $(1-\beta_{t+1})\nabla U({\mathbf{x}}_{t})$, we obtain
${\mathbf{z}}_{t+2}-$ $\nabla U({\mathbf{x}}_{t+1})=(1-\beta_{t+1})({\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t}))$
$-(1-\beta_{t+1})(\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t}))+\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})$ (68)
where observe that ${\mathbb{E}}_{{\boldsymbol{\xi}}_{t}}[\nabla U({\mathbf{x}}_{t})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})]={\mathbb{E}}_{{\boldsymbol{\xi}}_{t}}[\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})]=0$. Therefore, we have that
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\frac{\left\|{\mathbf{z}}_{t+2}-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}]={\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\bigg{]}$
$~{}~{}~{}~{}+{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}(1-2\beta_{t+1})^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
$~{}~{}~{}-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}-\beta_{t+1}\big{(}\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1}))\big{\|}^{2}\Big{]}$ (69)
@@
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}(1-2\beta_{t+1})^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
$~{}~{}~{}~{}~{}-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}-\beta_{t+1}\big{(}\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1}))\big{\|}^{2}\Big{]}$
$\overset{(a)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\bigg{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{\beta_{t+1}^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\bigg{]}$
$\overset{(b)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\bigg{]}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{\beta_{t+1}^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\bigg{]}$ (70)
where in (a) we have used the inequality ${\mathbb{E}}[\left\|\mathsf{X}-{\mathbb{E}}[\mathsf{X}]+\mathsf{Y}\right\|^{2}]\leq 2{\mathbb{E}}[\left\|\mathsf{X}\right\|^{2}]+2{\mathbb{E}}[\left\|\mathsf{Y}\right\|^{2}]$ for any random variables $\mathsf{X}$ and $\mathsf{Y}$ with bounded variances. And in (b) we have used ${\mathbb{E}}[\left\|\mathsf{X}-{\mathbb{E}}\left[\mathsf{X}\right]\right\|^{2}]\leq{\mathbb{E}}[\left\|\mathsf{X}\right\|^{2}]$. Dividing both sides by $\tfrac{1}{16L^{2}}$, taking full expectation in (69) and using $\beta_{t+1}\leq\frac{1}{2}$ (27) can be rewritten as,
$\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\right]\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]$
$~{}~{}+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}(1-2\beta_{t+1})}\right]$ (71)
@@
$\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t+1}\right\|^{2}}{\eta_{t}(1-2\beta_{t+1})}\right]$
$\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\frac{\left\|{\mathbf{x}}_{t+1}-{\mathbf{x}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}(1-2\beta_{t+1})}\right]$ (72)
$\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}(1-2\beta_{t+1})}\right].$ (73)

### v3

@@ IV Complexity analysis
Under a non-adaptive step-size setting with $\eta=\frac{{\bar{k}}}{T^{1/3}}$ and $\beta=c\eta^{2}$, we state the following corollary, which follows along the lines of Theorem 2 and is detailed in Appendix E.
Corollary 3.
Under Assumptions A1-A2 and A4-A10, along with conditions of Corollary 2 and Lemma 4, there exists some $1\leq t\leq{\mathcal{O}}(\epsilon^{-3/2})$ and ${\hat{{\boldsymbol{\lambda}}}}_{t}\in{\mathbb{R}}^{I}_{+}$, ${\hat{{\boldsymbol{\nu}}}}_{t}\in{\mathbb{R}}^{J}_{+}$ such that the point $({{\hat{{\mathbf{x}}}}_{t}},{\hat{{\boldsymbol{\lambda}}}}_{t},{\hat{{\boldsymbol{\nu}}}}_{t})$ is $\epsilon$-KKT optimal.
Under the adaptive step-size assumption, the SFO complexity bound in Theorem 2 matches that of STORM [20] and nearly matches Pstorm [33], STORM+ [30] achieving a complexity of $\mathcal{O}(\epsilon^{-3/2})$ for the unconstrained case. Notably, while STORM and Pstorm were proposed for unconstrained and convex-constrained problems, respectively, CoSTA extends this to non-convex constrained framework with a comparable SFO complexity. Furthermore, under the non-adaptive step-size setting, CoSTA attains the optimal convergence rate established by [14]. A key distinction of CoSTA, however, lies in its ability to achieve a near-optimal rate (i.e., ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$) under adaptive step sizes, a feature that distinguishes it from [14], which employs non-adaptive step sizes. It is remarked that the non-asymptotic results in Theorem 2 and Corollary 3 can also be used to infer the asymptotic convergence of the corresponding quantities.
@@ V-A Sparse Binary Classification
The problem of finding a sparse binary classifier can be written in the form of ($\mathcal{P^{\prime}}$) with the objective being the logistic loss function, i.e., $f({\mathbf{x}},\{{\mathbf{a}},b\})=\log(1+e^{-b{\mathbf{a}}^{\mathsf{T}}{\mathbf{x}}})$ and the constraint $g({\mathbf{x}})$ taking form of a smoothed sparsity inducing function, such as the $\ell_{1}$ norm or the minimax concave function [9]. The random variable here signifies the independent observations $\{{\mathbf{a}},b\}$. For our case we will take $\bar{g}$ to be minmax concave penalty (MCP) [9], given by:
$\bar{g}({\mathbf{x}})$ $=\lambda\|{\mathbf{x}}\|_{1}-\sum_{k=1}^{n}\bar{h}_{\lambda,\theta}([{\mathbf{x}}]_{k})$ (65)
where $[{\mathbf{x}}]_{k}$ denotes the $k$-th entry of ${\mathbf{x}}$ and $\bar{h}_{\lambda,\theta}$ is defined as:
$\bar{h}_{\lambda,\theta}(x)=\begin{cases}\frac{x^{2}}{2\theta},&|x|\leq\theta\lambda\\ \lambda|x|-\frac{\theta\lambda^{2}}{2},&|x|>\theta\lambda.\\ \end{cases}$ (66)
It can be noted that, even though $\bar{g}$ is non-smooth, a smoothed version of the MCP can be obtained by approximating $|x|$ as $\sqrt{x^{2}+\varrho}-\sqrt{\varrho}$, where $\varrho$ is a positive constant. Hence, instead of (65) with (66), a smoothed constraint can be obtained,
$g({\mathbf{x}})$ $=\sum\nolimits_{k=1}^{n}\Big{(}\lambda\sqrt{[{\mathbf{x}}]^{2}_{k}+\varrho}-\lambda\sqrt{\varrho}-h_{\lambda,\theta}([{\mathbf{x}}]_{k})\Big{)}$ (67)
@@
where $g$ is smoothed version of $\bar{g}$ and $h$ is the smoothed version of of $\bar{h}$. Since the objective is already convex, its running surrogate ${\tilde{f}}$ can be constructed using ${\hat{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t}),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ and (6) as
@@
${\tilde{f}}({\mathbf{x}},{\mathbf{x}}_{t},{\mathbf{z}}_{t},{\boldsymbol{\xi}}_{t})=f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+\langle\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t})+(1-\beta_{t})({\mathbf{z}}_{t}-\nabla f({\mathbf{x}}_{t-1},{\boldsymbol{\xi}}_{t})),{\mathbf{x}}-{\mathbf{x}}_{t}\rangle+\frac{\mu}{2}\left\|{\mathbf{x}}-{\mathbf{x}}_{t}\right\|^{2}$ (68)
where ${\boldsymbol{\xi}}_{t}=\{{\mathbf{a}}_{t},b_{t}\}$. On the other hand, the surrogate of $h$ can be constructed as,
$\tilde{h}_{\lambda,\theta}(x,x_{t})$ $=h_{\lambda,\theta}(x_{t})+\nabla h_{\lambda,\theta}(x_{t})(x-x_{t})$ (69)
=$=\begin{cases}\frac{x_{t}x}{\theta}-\frac{x_{t}^{2}}{2\theta},&|x_{t}|\leq\theta\lambda\\ \frac{\lambda\varrho}{\sqrt{x_{t}^{2}+\varrho}}-\lambda\sqrt{\varrho}-\frac{\theta\lambda^{2}}{2}+\frac{\lambda x_{t}x}{\sqrt{x_{t}^{2}+\varrho}},&|x_{t}|>\theta\lambda.\\ \end{cases}$ (70)
Hence, the surrogate of the smoothed constraint (67) at point ${\mathbf{x}}_{t}$ becomes
@@
${\tilde{g}}({\mathbf{x}},{\mathbf{x}}_{t})$ $=\sum\nolimits_{k=1}^{n}\Big{(}\lambda\sqrt{[{\mathbf{x}}]^{2}_{k}+\varrho}-\tilde{h}_{\lambda,\theta}([{\mathbf{x}}]_{k},[{\mathbf{x}}_{t}]_{k})-\lambda\sqrt{\varrho}+\tfrac{L}{2}([{\mathbf{x}}]_{k}-[{\mathbf{x}}_{t}]_{k})^{2}\Big{)}$ (71)
@@
It is worthy to note that 71 fulfills A9. We compare the performance of our algorithm with that of the LCPP method, which was proposed in [9] for solving stochastic non-convex optimization problems with sparsity inducing non-convex constraints. It is proximal point algorithm that solves the a sequence of convex subproblems with gradually relaxed constraint levels.
@@
For our simulations use two dfatasets: MNIST and gisette, details of which are summarized in Table II. For the MNIST dataset, we consider the binary classification problem of classifying the handwritten digit 5 from the other digits. As the dataset consists of 10 classes this can be achieved by assigning a label $+1$ to digit $5$ and $-1$ to other digits. Likewise, gisette dataset also consists handwritten digits, but here the classification is between the often confused digits 4 and 9.
@@
TABLE IV: Average time per iteration in seconds for $\sigma$ = 0.2, $T_{f}$ = 15, environmental configuration 4, and different values of $T$.
@@ VI Conclusion
We have considered the problem of minimizing a stochastic non-convex optimization problem with non-convex constraints. While the problem has been well-studied, the state-of-the-art algorithms have suboptimal oracle complexity of ${\mathcal{O}}(\epsilon^{-2})$. The recently proposed algorithm in [14] achieves the optimal rate of ${\mathcal{O}}(\epsilon^{-3/2})$, but this result is only valid under the condition that the initial point is nearly feasible with non-adaptive step sizes. In contrast, we propose the CoSTA algorithm, which introduces recursive momentum updates to achieve an almost optimal oracle complexity of ${\tilde{\mathcal{O}}}(\epsilon^{-3/2})$ under adaptive step size, and achieves an optimal oracle complexity with non-adaptive step size. The proposed algorithm is developed within the SCA framework, and requires convex surrogates of the objective and constraint functions at every iteration. We also introduced a parameterized version of the Mangasarian-Fromovitz constraint qualification that allows us to ensure that the intermediate subproblems are strictly feasible and the intermediate dual variable stay bounded, resulting in an oracle complexity bound that depends only on the problem parameters. Finally, the empirical performance of the proposed CoSTA algorithm is compared with the other state-of-the-art algorithms across two different applications: energy optimal trajectory design and sparse binary classification. The proposed algorithm outperforms the state-of-the-art algorithms on the energy optimal trajectory design problem and performs at par with the LCPP algorithm on the sparse binary classification problem.
@@ Appendix A Proof of Lemma 1
The proof follows along the lines of [20, Lemma 2]. Expanding ${\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})$ from (7) and introducing $(1-\beta_{t+1})\nabla U({\mathbf{x}}_{t})$, we obtain
[…]
$+\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})$ (74)
where observe that ${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\nabla U({\mathbf{x}}_{t})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})]={\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})]=0$. Therefore, we have that
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}[\tfrac{\left\|{\mathbf{z}}_{t+2}-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}]$ $={\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|{\mathbf{z}}_{t+1}-\nabla U({\mathbf{x}}_{t})\right\|^{2}}{\eta_{t}}\bigg{]}+{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})$
[…]
@@
${\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\Big{[}\eta_{t}^{-1}\big{\|}(1-\beta_{t+1})\big{(}\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t})+\nabla U({\mathbf{x}}_{t+1})-\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})\big{)}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\beta_{t+1}\big{(}\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1}))\big{\|}^{2}\Big{]}$
$\overset{(a)}{\leq}2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{(1-\beta_{t+1})^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla f({\mathbf{x}}_{t},{\boldsymbol{\xi}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}+2{\mathbb{E}}_{{\boldsymbol{\xi}}_{t+1}}\bigg{[}\frac{\beta_{t+1}^{2}\left\|\nabla f({\mathbf{x}}_{t+1},{\boldsymbol{\xi}}_{t+1})-\nabla U({\mathbf{x}}_{t+1})\right\|^{2}}{\eta_{t}}\bigg{]}$
[…]
where in (a) we have used the inequality ${\mathbb{E}}[\left\|\mathsf{X}-{\mathbb{E}}[\mathsf{X}]+\mathsf{Y}\right\|^{2}]\leq 2{\mathbb{E}}[\left\|\mathsf{X}\right\|^{2}]+2{\mathbb{E}}[\left\|\mathsf{Y}\right\|^{2}]$ for any random variables $\mathsf{X}$ and $\mathsf{Y}$ with bounded variances. And in (b) we have used ${\mathbb{E}}[\left\|\mathsf{X}-{\mathbb{E}}\left[\mathsf{X}\right]\right\|^{2}]\leq{\mathbb{E}}[\left\|\mathsf{X}\right\|^{2}]$. Dividing both sides by $\tfrac{1}{16L^{2}}$, taking full expectation in  (75) and using $\beta_{t+1}\leq\frac{1}{4}$, (75) can be rewritten as,
[…]
$~{}~{}~{}~{}+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}}\right]$ (77)
@@
[…]
$\leq\frac{1}{16L^{2}}{\mathbb{E}}\left[\frac{(1-\beta_{t+1})\left\|{{\boldsymbol{\varepsilon}}}_{t}\right\|^{2}}{\eta_{t}}\right]+\frac{1}{8}{\mathbb{E}}\left[\eta_{t}\left\|{\boldsymbol{\delta}}_{t}\right\|^{2}\right]+\frac{1}{8L^{2}}{\mathbb{E}}\left[\frac{\beta_{t+1}^{2}G_{t+1}^{2}}{\eta_{t}}\right].$ (79)